# A Compact, Low-cost Force and Torque Sensor for Robot Fingers with LED-based Displacement Sensing

### v1

@@ VI-A Data Collection
We selected 10 contact locations, spaced equally around the circumference of the finger. At each location, we collected data for 15 different contacts. Every set of three contacts was collected at different heights between the top, middle, and bottom of the finger, and focused on a different force range, to ensure full coverage. Forces were grouped into below 1N, around 1N, and around 2N, with 3 contacts for each force grouping followed by 6 more contacts split randomly between them. Contacts were done by pushing against the finger for between 2 and 5 seconds. For the test set, this process was repeated, but with only 1 contact per surface for each force grouping. The test set was randomly split across the height groupings to be representative of the different torque-force pairings.
During data collection, the sensor was measured sending data slightly upwards of 2.5KHz, but data was collected at 500Hz due to the limitations of the Virtual Machines used for data collection. We also observed that, for this version of the sensor, 4 LED signals exhibited high noise level, likely an artifact of traces along the PCB. The signals were thus dropped from the dataset during training. This left the model with 20 signals to learn from rather than the full 24. Resolving the electrical issues that caused these noisy signals is one of the areas for improvement.
VI-B Learning Algorithms
Before training the model, we filtered our data with a 45 Width Median Filter to reduce noise. The ground truth data was filtered via a 13 width Median Filter. Given our rate of 500Hz, this introduces under a tenth of a second delay. Additionally, we threshold all data points with a total force magnitude below .1N as ”no contact” when training our model. To prevent dataset over saturation of ”no contact” points, we down-sampled these points so they only comprise 10% of the training data. Additionally, rather than training directly on the raw signals, we trained the model on the change in signals, with respect to a no-contact baseline taken as the mean of 50 points from the start of the each data file. Finally, both features and labels are normalized around a mean of 0 and a standard deviation of 1.
For all experiments on the sensor, we use feed forward neural network architectures. The model consists of 6 layers with ReLU activation functions, broken up evenly between shared layers and 6 output heads, 1 for each of the 6 Force and Torque DOFs. The 3 shared layers were all size 128, while the output heads layers were 64, 32, then 1. The final layer does not have an activation function. We use mean squared error as a loss function. The network is trained for 50 epochs with ADAMoptimizer, a batch size of 2000 and an initial learning rate of 0.001. Additionally, the model reported on here uses time series data. Data from the four most recent time-steps are concatenated into a single linear input before being fed into the model.
VI-C Results and Discussion
The overall results of our sensor are shown in Table II and Figure 7. In particular, Table II show the performance for each direction of force and torque, as well as overal force and torque magnitude. Figure 7 plots mean absolute error as a function of applied force/torque. In addition, Figure 6 illustrates the performance of the sensor in detail using all six force/torque dimensions over two contacts, different in terms of the quality of the model fit.
We observe that the sensor provides accurate measurements in both force and torque for the x and y axes, but less so for the z axis. Noise levels are generally low, and predictions are accurate in terms of overall force magnitude, but can occasionally suffer from accuracy at higher forces in individual DOFs. We note again that the test data consisted of contacts never seen in training. The model was evaluated primarily on the 0-2N range of forces, but experimentation showed that training on data that extended beyond this range proved useful for the models learning.
@@ VI-C Results and Discussion
TABLE II: Accuracy Reporting for the F/T Model. X, Y, and magnitude results are quite promising, but Z still requires further data collection and training before conclusions can be drawn. Despite the absolute error being in the same threshold, the magnitude of Z forces is much smaller leading to a higher percentage error. This is most evident in the R2 results which shows the much higher correlation for x and y then for z
@@
We believe that one likely source of error is hysteresis introduced by the PDMS. This is particularly prevalent with larger forces, where signals will take some time to settle back to their baselines, which introduces noise to forces done before the resettling. It also introduces noise into the zero-force data. The physical sensors also introduced light bleed and reflections that were not present in our test set ups. With the wider emission cones caused by the PDMS, signals from emitters which were supposed to be picked up only by the cluster directly opposite of them were bleeding onto the other clusters. Additionally, the emitters were reflecting back onto the board they were on, creating signal even if only one board was powered. We believe that future versions of our sensor could leverage this effect to help increase z-sensitivity as we train more deliberately for Fz and Tz in the future.
We also verified our assumption that noisy raw LED signals are harmful to learning, by trained models with and without the 4 such signals observed to exhibit high noise. We found that the model without any of the 4 most noisy signals learned best. This leaves us optimistic that developing a version with the noise reduced via improved traces will lead to better predictions. Future versions will also analyze the possibility of removing some of the LEDs to shrink the overall form factor.
VII Conclusion and Future Work
@@
In this paper, we introduce a new method for building low-cost, compliant 6-axis F/T sensors designed to meet the needs for integration into robot hands. Our method relies on sensing the displacement between two sensor plates, each mounted with LEDs acting as light emitter or receivers. We analyzed such an LED-LED pair in isolation, and found it to be highly sensitive to relative displacements of the emitter and the receiver. We thus designed a sensor with 6 LEDs acting as emitters and 24 LEDs acting as receivers, distributed between the top and bottom plates of the sensor.
@@
We showed that a simple supervised learning model trained on data from the 24 receiver LEDs can learn to map changes in these signals to external forces and torques applied between the two plates via the induced displacements. The sensor output was particularly accurate for force/torques along/around the x and y axes, with additional work needed to improve sensitivity for the z axis.
@@
Future work of the sensor falls under two categories: further improvement of the sensor, and implementation within actual use cases. Sensor improvement in turn will have mechanical, electronic, and learning components. Mechanically, we would like to increase the overload protection, reduce hysteresis, and ensure more consistency in the parts for better alignment. Electronically, we will continue to look for sources of noise in LED signals. We also want to minimize the form factor as much as possible, so further experiments would include use of fewer LEDs in each cluster (or potentially even fewer clusters). Learning wise, we will further tune the model as we gather more data for Z axis and explore alternate architectures with both time series and single-snapshot models.
For deployment of the sensor, we look to take advantage of the features of the sensor which make it particularly attractive for integration in robot hands. In particular, the form factor and ease of integration (the sensor directly connects to a microcontroller board with no additional amplification electronics) are highly appealing. Figure 8 shows a possible design of an anthropomorphic hand with one of our sensors at the base of the distal phalanx of each finger. Such integration will also allow us to test the hypothesis that slight compliance in the sensing element can prove beneficial, or at least non-detrimental, to dexterous manipulation.
As motor learning for dexterous manipulation makes continuous advances, we believe that sensing will prove key for future performance of robot hands. Having a multitude of sensors at our disposal, each with various strengths, will increase the ability of combining the right sensor with the right motor learning method to achieve novel capabilities. We hope that the sensing method presented here represents a step in this direction.

### v2

@@ V-A Data Collection
We selected 10 contact locations, distributed evenly around the circumference of the finger. Contacts were applied by pressing against the printed finger with our fingertip for 2 and 5 seconds. For the training set, we collected data for 15 different contacts at each location. These contacts were grouped into sets of three, each recorded at different heights—top, middle, and bottom of the finger. Each set focused on a specific force range to ensure comprehensive coverage. The test sets are collected with the same procedure, but included only 1 contact per surface for each contact position. The test set was randomly split across the height groupings to be representative of the different torque-force pairings.
During data collection, the sensor operated at a sampling rate slightly above 2.5 kHz. However, data was down-sampled and recorded at 250 Hz due to the practical requirements of robotic manipulation tasks.
V-B Data pre-processing
Prior to training, we applied a series of pre-processing techniques to improve data quality and model performance. To reduce high-frequency noise, we used an 11-point median filter on the raw sensor readings and a 5-point median filter on the ground truth data. Given our data collection rate of 250 Hz, this filtering process introduces a delay of less than 0.1 seconds, which is negligible for our application.
To distinguish between contact and non-contact states, we thresholded all data points with a total force magnitude below 0.1 N as “no contact.” However, to prevent the dataset from being overwhelmingly dominated by these points, we downsampled them to ensure they constitute only 10% of the training data.
Additionally, instead of training directly on absolute raw signals, we extracted features based on their relative changes from a no-contact baseline. This baseline was defined as the mean of the first 50 data points in each recorded dataset. To further stabilize training and maintain numerical consistency, we normalized all features and labels prior to training. These pre-processing steps help the model focus on signal variations rather than absolute values, enhancing generalization.
V-C Learning Algorithms
For all experiments on the sensor, we employ an updated ResNet-based architecture inspired by He et al. [26]. Our model consists of a shared feature extraction backbone followed by six independent task-specific heads, one for each of the six DOFs in force and torque prediction.
@@ V-C Learning Algorithms
TABLE II: Accuracy Reporting for the F/T Model. X, Y, and magnitude results are quite promising, but Z still requires further data collection and training before conclusions can be drawn. Despite the absolute error being in the same threshold, the magnitude of Z forces is much smaller leading to a higher percentage error. This is most evident in the R2 results which shows the much higher correlation for x and y than for z
The key component of our model is the residual block with 1D convolutional layers. Each residual block integrates 1D convolutional layers, skip connections and ReLU activation functions to enhance feature extraction while mitigating vanishing gradients.
The input to the model is a 24-channel time-series sequence, corresponding to the preprocessed sensor signals. A 1D convolutional layer with an enlarged kernel is applied before the residual blocks to capture global patterns in the sensor signals. This is followed by a shared backbone, which is a 1D ResNet specifically designed for sequential sensor data. The backbone consists of multiple residual blocks that efficiently capture spatial and temporal dependencies.
After the shared backbone, the model branches into six independent prediction heads, each dedicated to estimating a specific force or torque component. Each head consists of additional residual layers, a global average pooling layer, and a fully connected output layer that generates a single scalar value. This multi-task learning framework enables specialized feature extraction while leveraging shared representations, thereby enhancing generalization across different force and torque components.
The model is trained using a mean squared error (MSE) loss function, applied separately to each output head. Training is conducted for 100 epochs using the Adam optimizer with a batch size of 2048 and an initial learning rate of 1e-3, which is adjusted dynamically using a LambdaLR scheduler.
V-D Results and Discussion
The overall results of our sensor’s contact prediction are presented in Table II and Figure 8. In particular, Table II details the performance for each force and torque direction, as well as the overall force and torque magnitudes. Figure 8 visualizes the mean absolute error as a function of applied force and torque. In addition, Figure 7 provides a detailed performance analysis, showcasing all six force/torque dimensions across three exemplar contacts, each differing in the contact positions. We note again that the test data consisted of contacts never seen in training.
We observe that the sensor provides accurate measurements in both force and torque for the x and y axes, but less so for the z axis. Noise levels are generally low, and predictions are accurate in terms of overall force magnitude, but can occasionally suffer from accuracy at higher forces in individual DOFs. The model was evaluated primarily on the 0-5 N range of forces, but experimentation showed that training on data that extended beyond this range proved useful for the models learning.
@@ V-D Results and Discussion
We believe that one likely source of error is hysteresis introduced by the PDMS and quantified in Sec. IV. This is particularly prevalent with larger forces, where signals will take some time to settle back to their baselines, which introduces noise to forces done before the resettling. It also introduces noise into the zero-force data.
VI Conclusion and Future Work
@@
In this paper, we introduce a low-cost, compliant sensor designed to be easily integrated in a robot finger and provide information on the displacement induced by contact with the world. Our method relies on sensing the displacement between two rigid plates, each mounted with LED emitters and receivers. We analyzed such an LED-LED pair in isolation, and found it to be highly sensitive to relative displacements of the emitter and the receiver. We thus designed a sensor with six LEDs acting as emitters and 24 LEDs acting as receivers, distributed between the top and bottom plates of the sensor.
@@
To characterize the information contained in the raw signals from our sensor, we tested wether it is possible to reconstruct overall finger wrench information, using a commercial force/torque sensor ars ground truth. Our results showed that a simple supervised learning model trained on data from the 24 receiver LEDs can learn to map changes in these signals due to displacement of the flexure to external forces and torques applied between the two plates. The sensor output was particularly accurate for force/torques along/around the x and y axes, with additional work needed to improve sensitivity for the z axis.
Although our sensor does not outperform the accuracy or precision of commodity F/T sensors, we believe that it offers useful information - especially along/about the axes most prevalent for manipulation - in a compact, robust and low-cost package. Modern learning-based manipulation methods often operate directly on raw, even uncalibrated sensor data [27, 28], thus converting to an intermediate representation such as the one used here for sensor characterization may be unnecessary in the future. Considering the strong SNR of the raw signals of our sensor, we believe that it can be used in a similar manner for policy learning for manipulation.
@@
Future work of the sensor falls under two categories: further improvement of the sensor, and implementation within actual use robot hands. Sensor improvement in turn will have mechanical, electronic, and learning components. Mechanically, we would like to reduce hysteresis, simplify the manufacturing, and improve the consistency between sensors. We would also like to reduce the stiffness in the z-axis so that we get more sensitivity in that direction. Electronically, we will continue to look for sources of noise in LED signals. We also want to minimize the form factor as much as possible, so further experiments would include use of fewer LEDs in each cluster (or potentially even fewer clusters). Learning wise, we will further tune the model as we gather more data for Z axis and explore alternate architectures with both time series and single-snapshot models.
@@
We look to take advantage of the features of the sensor which make it particularly attractive for integration in robot hands. In particular, the form factor and ease of integration (the sensor directly connects to a micro-controller board with no additional amplification electronics) are highly appealing. Figure 9 shows a possible design of an anthropomorphic hand with one of our sensors at the base of the distal phalanx of each finger. Such integration will also allow us to test the hypothesis that slight compliance in the sensing element can prove beneficial, or at least non-detrimental, to dexterous manipulation.
As motor learning for dexterous manipulation makes continuous advances, we believe that sensing will prove key for future performance of robot hands. Having a multitude of sensors at our disposal, each with various strengths, will increase the ability of combining the right sensor with the right motor learning method to achieve novel capabilities. We believe that the sensor presented in this paper represents a step in this direction.